A medication order calls for two tablets Q.D. the first day, one tablet b.i.d. the second and third days and one tablet q.d. thereafter for a total of two weeks. How many tablets must be dispensed in total?
Answer #1

A medication order calls for two tablets Q.D. the first day=2x4= 8 tablets first day

Then one tablet b.i.d. on second day = 1x2= 2 tablets

And third day also one tablet b.i.d. = 1x2= 2 tablets

Then one tablet q.d thereafter for a total of two weeks= 1X1 (14-3) = 11

Total tablets must be dispensed= 8+2+2+11= 23 tablets answer